I do not understand where to even begin in a problem like this one. Please help me to understand how to reach the answer and why it is so. Thank you! :)
Let (f)x = –2x + 4 and g(x) = –6x – 7. Find f(x) – g(x)

Found 2 solutions by fcabanski, jsmallt9:
Answer by fcabanski(1391)

Plug in the given values for f(x) and g(x)


If f(x) = 1 and g(x) = 2 it would be 1-2=-1


f(x) - g(x) = -2x + 4 -(-6x-7) Distribute that -


-2x + 4 + 6x + 7 = 4x + 11

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)


Answer by jsmallt9(3758)

Since (f)x = –2x + 4 and g(x) = –6x – 7 and you want to find f(x) - g(x), just replace f(x) with -2x+4 and g(x) with -6x-7:
f(x) - g(x)
(-2x+4)-(-6x-7)
[Important! Notice the use of parentheses! It is an extremely good habit to use parentheses with making substitutions (especially if you are substituting in multiple terms like we are here).]

Since subtractions cause problems (they are harder than additions and they are not commutative so you can't change the order), I am going to change the subtractions to additions of the opposite:
(-2x+4)-(-6x+(-7))
(-2x+4)+(6x+7) (Note how the signs changed for both terms in the parentheses that had the "-" in front!)
Since we now have all addition we can freely change the order and grouping of the terms. Reordering and grouping the like terms together we get:
(-2x+6x)+(4+7)
Adding the like terms:
4x+11

So f(x) - g(x) = 4x+11